# Taking the Garbage Out of Data-Driven Prediction Across Climate Timescales

### v1

@@ 3.2 Working with categorical inputs and outputs
Sometimes prediction problems require predicting a category or class (e.g., El Niño or La Niña). Several methods should be considered when working with categorical data. If the data have ordered ranks (e.g., labeling 2-m temperatures as below average, average, and above average), then integer encoding, where a unique integer value is assigned to the respective category (e.g., 0, 1, or 2), is appropriate. Classes representing unordered data (e.g., European weather regimes; Grams et al., 2017) should use one-hot encoding, whereby a single value in the vector is set to one (corresponding to the category), and the others are left as zero. For the seven European weather regimes, for example, a day labeled “Atlantic Ridge” may be encoded [0,0,1,0,0,0,0,0] (the eighth category representing “no regime”; Grams et al., 2017). One-hot encoding prevents the AI/ML model from interpreting the categories as having an ordinal relationship. Binary categorical problems (e.g., severe or non-severe thunderstorms) can choose between integer or one-hot encoding. Missing values can be assigned as a separate category/class (e.g., “missing”) before encoding, or if there are few missing values, they may be deleted.
Some climate prediction problems, particularly when working with extreme events, may have data with very few events and many “null” events – e.g., if classifying days with 2-m temperatures exceeding the 99th percentile, there will be 1 event for every 100 days, on average. In this example, the extreme heat days comprise a minority class compared to the non-extreme heat days. Hence, the user is faced with what is known as class imbalance, meaning there are very few “hits” for the AI/ML model on which to train, and thus, overall poor performance by the AL/ML model (Molina et al., 2023b). Since minority classes can be important in climate prediction studies, one should consider resampling techniques to address the imbalance. For example, the size of the minority class can be increased by randomly duplicating existing samples or using algorithms to create synthetic samples, known as data augmentation. One such data augmentation method is the Synthetic Minority Oversampling Technique (SMOTE; Chawla et al., 2002). Care should be taken when oversampling autocorrelated data; techniques for time series imputation such as time-sliced SMOTE can be used (Baumgartner et al., 2022). Undersampling of the majority class can also be employed to reduce the imbalance (e.g., Gensini et al., 2021; Rivera Tello et al., 2023). Class weights can also be applied, where larger magnitude weights can be assigned to the minority class for an added penalty in its erroneous classification during training. Importantly, the minority class should contain a diverse set of examples from which to learn. Thus, we recommend the input data to have a good $N_{\mathrm{eff}}$ as opposed to focusing on the total sample size.
@@ 4.1 Case Study #1: Subseasonal weather regimes
Weather regimes are large-scale, persistent, and recurrent atmospheric patterns useful for subseasonal predictions due to their relationship with surface weather anomalies (e.g., Molina et al., 2023a). These regimes are often defined using 500 hPa geopotential height (Z500) anomalies over a specified domain (e.g., North America) and $k$-means clustering. Here, we demonstrate how seemingly minor differences in preprocessing choices can lead to a different number of preferred weather regimes with different spatial characteristics. To highlight this sensitivity, we will calculate the North American weather regimes in three different ways, detailed below. Daily-mean Z500 comes from ERA5 (Hersbach et al., 2020) and is regridded to a 1° horizontal grid spacing using the nearest neighbor method. The analysis region is chosen as 20°N–80°N, 180°–30°W, as in Lee et al. (2023). A daily climatology spanning 1940–2023 is established, utilizing a 60-day centered rolling window method. A 10-day lowpass filter is applied to the anomalies, and they are also detrended by subtracting a latitudinal-weighted third-degree polynomial fit for each day of the year. The data are then standardized using the 60-day centered running mean of the area-averaged Z500 anomaly standard deviation for each day of the year. Subsequent steps are detailed in Pérez-Carrasquilla and Molina (2025).
We conduct three experiments (Table 2) for this case study. The “control” experiment aligns with the definitions and methods used by Lee et al. (2023) except for using the 1940–-2023 climatological base period. “Experiment 1” involves standardizing the detrended anomalies using grid-cell daily-averaged standard deviation instead of using the domain-averaged standard deviation, as done in Lee et al. (2023). “Experiment 2” involves using the climatological period 1979–-2023. All other preprocessing steps are consistent among the experiments. The optimal number of clusters for each experiment is found using the intercluster correlation, which is the correlation between the centroid coordinates. An intercluster correlation at zero or just below zero is preferred. The optimal number can vary depending on the chosen metric and more than one metric can be used for robustness; however, here we focus on sensitivity to preprocessing choices, not the chosen metric(s).
@@
Control Lee et al. (2023), but with 1940–-2023 climatology and 3rd-degree polynomial detrending. 6 Alaskan Low (AL), Alaskan Ridge (AR), North Atlantic High (NAH), Pacific Trough (PT), Pacific Ridge (PR), and Greenland High (GH)
@@
Figure 3 summarizes the results of the experiments. The control experiment results in the six “preferred” North American weather regimes (Fig. 3a). In contrast, Experiment 1 yields four regimes, whereas Experiment 2 results in six preferred regimes (Fig. 3a). In Experiment 1, local standardization alters the spatial patterns of the Greenland High, Pacific Ridge, and Pacific Trough regimes (Figs. 3e, g, h). Local standardization also leads to the emergence of the Central US High regime (not shown), which was absent in the control experiment. In Experiment 2, a shorter climatology eliminates the Alaskan Low regime, suggesting its frequency may have diminished in recent decades (Fig. 3b). The decision to use grid-cell rather than regional or global standardization stems from the necessity to capture localized signals, potentially useful for subseasonal high-impact weather events. A shorter climatology may be preferred if polynomial detrending is ineffective in removing regional trends. The methodology used in the control experiment is favored due to the need to capture large-scale patterns (regional standardization) and the intention to investigate trends in weather regimes (longer climatology).
@@ 4.2 Case Study #2: Predicting temperature anomalies in the Southwest US
Time series prediction is commonly performed in climate science, particularly for understanding and predicting climate patterns. Furthermore, many climate modes and large-scale processes are captured through time series indices, such as the El Niño-Southern Oscillation (ENSO). Forecasting time series evolutions can aid in weather forecasting and prediction of extreme events, such as heat waves. To illustrate the impact on prediction skill from common preprocessing missteps, we evaluate a simple neural network regression model to predict the monthly temperature anomaly in the Southwest US. Note that the neural network built for this case study is intended to illuminate various preprocessing effects on outcomes and not for actual prediction. For many applications, the observational record is often too short for developing such neural network models, as time series can exhibit strong temporal autocorrelation. Many weather and climate forecasting applications are developed using multi-global climate model data (e.g., Ham et al., 2019; Rivera Tello et al., 2023), which increases $N_{\mathrm{eff}}$ by at least an order of magnitude. However, we have ensured that our model is not overfit to the training data and is sufficient in demonstrating preprocessing differences.
@@
The case study is set up as follows. Monthly-mean average temperatures are taken from the Berkeley Earth Surface Temperatures dataset (Rohde and Hausfather, 2020) from 1900–2025 and averaged across the Southwest US (29°N-39°N; 104°W-117°W). We input the current temperature anomaly and three lagged time steps of the timeseries to the neural network (3 layers, 10 nodes each) to predict the Southwest US temperature anomaly 4 months later. We focus on three specific preprocessing components: 1) the data split, 2) the climatological period, and 3) the detrending period. We construct several variants of the model (i.e., experiments), highlighting the use and misuse of each of these three preprocessing steps (Table 3). The clean preprocessing experiment represents the case in which the recommended preprocessing steps highlighted in this article are fully followed. The data split between the training and validation and validation and testing periods is an 18-month gap to reduce data leakage. Additionally, the climatology is computed over a 30-year period from the middle of the training period from January 1941 to December 1970. The linear trend (0.0289°C decade-1) is also computed over the full training period of January 1900 to December 1979. Skill for all experiments is computed via mean absolute error (MAE) and mean squared error (MSE) with lower MAE/MSE indicating higher skill.
@@
  • For the trend experiment (Fig. 4, orange line), the linear trend was computed over the full dataset (0.1048°C decade-1; January 1900 to December 2024). The trend over the full data set is nearly 3 times higher than the trend over the training period. Thus, the error is lower for the trend predictions, due to knowledge of the increased trend during model training. However, this trend from the testing period would not be known in a true testing sense where the testing data are completely unseen. Thus, the model’s performance is inflated; the trend has the largest impact on the overall skill in this case study.
@@
  • Finally, for the split-trend-climo experiment (Fig. 4, purple line), we combine the preprocessing missteps from the three previous experiments. The result is a slightly lower error than those three experiments, again producing artificially inflated skill.
Figure 4: Temperature anomaly (black; the “truth”) and neural network predictions using test datasets preprocessed in different ways: (1) trend computed during the test period (trend; orange), (2) climatology computed during the validation and test periods (climo; green), (3) data splits with potential leakage (split; red), and (4) a combination of the trend, climo, and split preprocessing steps (split_trend_climo; purple) (5) no data leakage (clean; blue). Corresponding skill scores are shown in parentheses: MAE (mean absolute error; °C) and MSE (mean squared error; °C2).
@@ 5 Conclusions
The use of AI/ML in climate prediction is rapidly expanding, introducing challenges with model design, skill assessment, and ultimately trust. A useful step towards building that trust is transparency in the process, including the initial problem design and data preprocessing. This work presents recommended steps for proper dataset preprocessing for different climate prediction problems. Such steps are recommended across applications and will serve as a way to make conscious choices when framing a prediction problem for climate timescales. The two case studies presented illustrate the importance of our recommended preprocessing steps and show how they can affect interpretation of the predictions. Understanding the importance of these preprocessing steps will likely lessen the frequent critique of the “black box” nature of AI/ML (McGovern et al., 2019). Hence, following the recommendations laid out in this article will move the community toward AI/ML applications for climate prediction that are transparent and fairly evaluated.
@@ Acknowledgements.
The authors would like to thank the US Climate Variability and Predictability Program (CLIVAR) for its support for this paper. M. J. M. was supported by the National Science Foundation (NSF) Grant #2425735. M. C. A. was funded, in part, by the National Oceanic and Atmospheric Administration (NOAA) Grant #NA24OARX431C0065-T1-01 and by the US Department of Energy’s (DOE) Office of Biological and Environmental Research (BER) Regional and Global Model Analysis (RGMA) program, as part of the Program for Climate Model Diagnosis and Intercomparison (PMCDI) Project. J. S. P. C. was supported by a University of Maryland Grand Challenges Seed Grant. T. B. received support from from the Horizon Europe project “Artificial Intelligence for enhanced representation of processes and extremes in Earth System Models (AI4PEX)” (Grant agreement ID: 101137682), funded by the Swiss State Secretariat for Education, Research and Innovation (SERI, Grant No. 23.00546). The scientific results and conclusions, as well as any view or opinions expressed herein, are those of the authors and do not necessarily reflect the views of NWS, NOAA, or the Department of Commerce.

### v2

@@ 3.2 Working with categorical inputs and outputs
Sometimes prediction problems require predicting a category or class (e.g., El Niño or La Niña). Several methods should be considered when working with categorical data. If the data have ordered ranks (e.g., labeling 2-m temperatures as below average, average, and above average), then integer encoding, where a unique integer value is assigned to the respective category (e.g., 0, 1, or 2), is appropriate. Classes representing unordered data (e.g., European weather regimes; grams2017) should use one-hot encoding, whereby a single value in the vector is set to one (corresponding to the category), and the others are left as zero. For the seven European weather regimes, for example, a day labeled “Atlantic Ridge” may be encoded [0,0,1,0,0,0,0,0] (the eighth category representing “no regime”; grams2017). One-hot encoding prevents the AI/ML model from interpreting the categories as having an ordinal relationship. Binary categorical problems (e.g., severe or non-severe thunderstorms) can choose between integer or one-hot encoding. Missing values can be assigned as a separate category/class (e.g., “missing”) before encoding, or if there are few missing values, they may be deleted.
Some climate prediction problems, particularly when working with extreme events, may have data with very few events and many “null” events – e.g., if classifying days with 2-m temperatures exceeding the 99th percentile, there will be 1 event for every 100 days, on average. In this example, the extreme heat days comprise a minority class compared to the non-extreme heat days. Hence, the user is faced with what is known as class imbalance, meaning there are very few “hits” for the AI/ML model on which to train, and thus, overall poor performance by the AL/ML model (molina2023review). Since minority classes can be important in climate prediction studies, one should consider resampling techniques to address the imbalance. For example, the size of the minority class can be increased by randomly duplicating existing samples or using algorithms to create synthetic samples, known as data augmentation. One such data augmentation method is the Synthetic Minority Oversampling Technique (SMOTE; chawla2002). Care should be taken when oversampling autocorrelated data; techniques for time series imputation such as time-sliced SMOTE can be used (baumgartner2022). Undersampling of the majority class can also be employed to reduce the imbalance (e.g., gensini2021; riveratello2023). Class weights can also be applied, where larger magnitude weights can be assigned to the minority class for an added penalty in its erroneous classification during training. Importantly, the minority class should contain a diverse set of examples from which to learn. Thus, we recommend the input data to have a good $N_{\mathrm{eff}}$ as opposed to focusing on the total sample size.
@@ 4.1 Case Study #1: Subseasonal weather regimes
Weather regimes are large-scale, persistent, and recurrent atmospheric patterns useful for subseasonal predictions due to their relationship with surface weather anomalies (e.g., molina2023). These regimes are often defined using 500 hPa geopotential height (Z500) anomalies over a specified domain (e.g., North America) and $k$-means clustering. Here, we demonstrate how seemingly minor differences in preprocessing choices can lead to a different number of preferred weather regimes with different spatial characteristics. To highlight this sensitivity, we will calculate the North American weather regimes in three different ways, detailed below. Daily-mean Z500 comes from ERA5 (hersbach2020) and is regridded to a 1° horizontal grid spacing using the nearest neighbor method. The analysis region is chosen as 20°N–80°N, 180°–30°W, as in lee2023. A daily climatology spanning 1940–2023 is established, utilizing a 60-day centered rolling window method. A 10-day lowpass filter is applied to the anomalies, and they are also detrended by subtracting a latitudinal-weighted third-degree polynomial fit for each day of the year. The data are then standardized using the 60-day centered running mean of the area-averaged Z500 anomaly standard deviation for each day of the year. Subsequent steps are detailed in perez-carrasquilla2025.
We conduct three experiments (Table 2) for this case study, each with 500 random centroid initializations to ensure robustness of resultant clusters. The “control” experiment aligns with the definitions and methods used by lee2023 except for using the 1940–-2023 climatological base period. “Experiment 1” involves standardizing the detrended anomalies using grid-cell daily-averaged standard deviation instead of using the domain-averaged standard deviation, as done in lee2023. “Experiment 2” involves using the climatological period 1979–-2023. All other preprocessing steps are consistent among the experiments. The optimal number of clusters for each experiment is found using the intercluster correlation, which is the correlation between the centroid coordinates. An intercluster correlation at zero or just below zero is preferred. The optimal number can vary depending on the chosen metric and more than one metric can be used for robustness; however, here we focus on sensitivity to preprocessing choices, not the chosen metric(s). Regime names were assigned subjectively based on similarities in Z500 anomalies, though distance or similarity metrics (e.g., Pearson correlation) can be used for more robust cluster alignment across experiments.
@@
Control lee2023, but with 1940–-2023 climatology and 3rd-degree polynomial detrending. 6 Alaskan Low (AL), Alaskan Ridge (AR), North Atlantic High (NAH), Pacific Trough (PT), Pacific Ridge (PR), and Greenland High (GH)
@@
Figure 3 summarizes the results of the experiments. The control experiment results in the six “preferred” North American weather regimes due to a relative “best” intercluster correlation value at $k=6$ compared to other $k$ values (Fig. 3a). In contrast, Experiment 1 yields four regimes, whereas Experiment 2 results in six preferred regimes (Fig. 3a). In Experiment 1, local standardization alters the spatial patterns of the Greenland High, Pacific Ridge, and Pacific Trough regimes (Figs. 3e, g, h). Local standardization also leads to the emergence of the Central US High regime (not shown), which was absent in the control experiment. In Experiment 2, a shorter climatology eliminates the Alaskan Low regime, suggesting its frequency may have diminished in recent decades (Fig. 3b). The decision to use grid-cell rather than regional or global standardization stems from the necessity to capture localized signals, potentially useful for subseasonal high-impact weather events. A shorter climatology may be preferred if polynomial detrending is ineffective in removing regional trends. A challenge with unsupervised learning is determining the “correct” final groupings, but the methodology used in the control experiment is preferred due to the need to capture large-scale patterns (regional standardization) and the intention to investigate trends in weather regimes (longer climatology).
@@ 4.2 Case Study #2: Predicting temperature anomalies in the Southwest US
Time series prediction is commonly performed in climate science, particularly for understanding and predicting climate patterns. Furthermore, many climate modes and large-scale processes are captured through time series indices, such as the El Niño-Southern Oscillation (ENSO). Forecasting time series evolutions can aid in weather forecasting and prediction of extreme events, such as heat waves. To illustrate the impact on prediction skill from common preprocessing missteps, we evaluate a simple neural network regression model to predict the monthly temperature anomaly in the Southwest U.S. Note that the neural network built for this case study is intended to illuminate various preprocessing effects on outcomes and not for actual prediction. For many applications, the observational record is often too short for developing such neural network models, as time series can exhibit strong temporal autocorrelation. Many weather and climate forecasting applications are developed using multi-global climate model data (e.g., ham2019; riveratello2023), which increases $N_{\mathrm{eff}}$ by at least an order of magnitude. However, we have ensured that our model is not overfit to the training data via implementing early stopping (e.g., patience=50, which indicates how many epochs to wait for a model’s performance to improve before stopping training) and is sufficient in demonstrating preprocessing differences.
@@
The case study is set up as follows. Monthly-mean average temperatures are taken from the Berkeley Earth Surface Temperatures dataset (rohde2020) from 1900–2025 and averaged across the Southwest US (29°N-39°N; 104°W-117°W). We input the current temperature anomaly and three lagged time steps of the timeseries to the neural network (3 layers, 100 nodes in the first layer, 50 nodes each in the second and third layer) to predict the Southwest US temperature anomaly 1 month later. We focus on three specific preprocessing components: 1) the data split, 2) the climatological period, and 3) the detrending period. We construct several variants of the model (i.e., experiments), highlighting the use and misuse of each of these three preprocessing steps (Table 3). The clean preprocessing experiment represents the case in which the recommended preprocessing steps highlighted in this article are fully followed. The data split between the training and validation and validation and testing periods is an 18-month gap to reduce data leakage. Additionally, the climatology is computed over a 30-year period from the middle of the training period from January 1941 to December 1970. The linear trend (0.0289°C decade-1) is also computed over the full training period of January 1900 to December 1979. Skill for all experiments is computed via mean absolute error (MAE) with lower MAE indicating higher skill. We first compute the MAE of the predictions using the test labels as they were originally (incorrectly) computed for each experiment, which results in inflated, “apparent” skill estimates. We then compute the adjusted MAE using the correctly cleaned test labels, which represents the true prediction error that would be observed in settings with no prior knowledge of the test data, such as real-time forecasting.
The model was trained using the Adam optimizer to minimize mean squared error loss. Additional parameters include a batch size of 64 and a learning rate of 0.00001. Parameters were selected to minimize loss for the clean experiment and the same parameters were used for all experiments.
@@
  • For the trend experiment (Fig. 4, orange line), the linear trend was computed over the full dataset (0.1048°C decade-1; January 1900 to December 2024). The trend over the full data set is nearly 3 times higher than the trend over the training period. Thus, the error is lower for the trend predictions, due to knowledge of the increased trend during model training. However, this trend from the testing period would not be known in a true testing sense where the testing data are completely unseen. Thus, the model’s performance is inflated, as shown by the adjusted MAE being much higher; the trend has the largest impact on the overall skill in this case study.
@@
  • Finally, for the split-trend-climo experiment (Fig. 4, purple line), we combine the preprocessing missteps from the three previous experiments. The result is a slightly lower error than those three experiments, resulting in artificially inflated skill when compared with the adjusted MAE.
Figure 4: Temperature anomaly (black; the “truth”) and neural network predictions using test datasets preprocessed in different ways: (1) trend computed during the test period (trend; orange), (2) climatology computed during the validation and test periods (climo; green), (3) data splits with potential leakage (split; red), and (4) a combination of the trend, climo, and split preprocessing steps (split_trend_climo; purple) (5) no data leakage (clean; blue). Corresponding skill scores are shown in parentheses. The MAE (mean absolute error; °C) is first calculated using the incorrectly computed test labels for each experiment, producing inflated skill estimates. The ”adjusted” MAE is calculated using the properly cleaned test labels to obtain an accurate measure of prediction error.
@@ 5 Conclusions
The use of AI/ML in climate prediction is rapidly expanding, introducing challenges with model design, skill assessment, and ultimately trust. A useful step towards building that trust is transparency in the process, including the initial problem design and data preprocessing. This work presents recommended steps for proper dataset preprocessing for different climate prediction problems. Such steps are recommended across applications and will serve as a way to make conscious choices when framing a prediction problem for climate timescales. The two case studies presented illustrate the importance of our recommended preprocessing steps and show how they can affect interpretation of the predictions. Understanding the importance of these preprocessing steps will likely lessen the frequent critique of the “black box” nature of AI/ML (mcgovern2019). Hence, following the recommendations laid out in this article will move the community toward AI/ML applications for climate prediction that are transparent and fairly evaluated.